# Sharp thresholds for spanning regular subgraphs

### v2

@@ 6.2. Improving fragments
For $F\in\mathcal{F}_{n}(\overrightarrow{D})$ that have an $(F,\mathbf{W})$-fragment of size at most $\ell_{0}+5\chi$, let $H=H(F)=F^{\prime}\cap F$ be a fragment from the set of all $(F,\mathbf{W})$-fragments of such a size such that $F^{\prime}$ is sampled according to a certain distribution $\mathcal{Q}_{F,\mathbf{W}}$ over $\mathcal{F}_{n}(\overrightarrow{D})$ that will be defined later in Section 6.3. We will need this specific distribution in order to prove that, on the last day, at least one fragment can be covered (see Section 6.7). In particular, this distribution has the following property: for every $(F,W)$ such that there is an $(F,W)$-fragment of size at most $\ell_{0}+5\chi$, we have that $\mathbb{P}(|\mathbf{F^{\prime}}\cap F|\leq\ell_{0}+5\chi)=1$, for $\mathbf{F}^{\prime}\sim\mathcal{Q}_{F,W}$.
@@
All graphs from $\mathcal{F}_{n}(\overrightarrow{D})$ contain every $D_{i}$, $i\in[\chi]$. Since $H=F\cap F^{\prime}$ for some $F^{\prime}\in\mathcal{F}_{n}(\overrightarrow{D})$, we get that every $D_{i}$ is a subgraph of $H$. Since the distance in $F$ between any two $D_{i}$ is bigger than $2C\sqrt{n}$, we get that the connected components of $H$ that contain different $D_{i}(F)$ are disjoint. Consider all inclusion-maximal closed subgraphs in $H$ with at least $10C\ln n/w$ vertices. For each such second power of a path $v_{1}\ldots v_{h}$, we “cut” it into pieces of length $\mu:=\left\lfloor 10C\ln n/w\right\rfloor$ in the following way: $v_{1}v_{2}$; $v_{3}\ldots v_{2+\mu}$; $v_{3+\mu}\ldots v_{2+2\mu}$; $\ldots$. We let the last piece $v_{3+j\mu}v_{4+j\mu}\ldots$ to be of length at least 2 and at most $\mu+1$. From every such inclusion-maximal closed subgraph, remove all pieces (i.e. all the listed vertices as well as the edges that contain them) but the first and the last one and glue the first and the last so that they make up the second power of a path $(v_{1}v_{2}v_{3+j\mu}v_{4+j\mu}\ldots)$. Let $\mathcal{P}_{H}=\{P_{1},\ldots,P_{s}\}$ be the set of all pieces of length $\mu$ that we have removed ($P_{1},\ldots,P_{s}$ respect the order induced by $F$). Clearly, the number of pieces $s$ is less than $32\ell_{0}/\mu<\chi/$ 3, where $\chi$ is defined in (32).
@@
We then choose different $j_{1},\ldots,j_{s}\in[\chi]$, according to some rule, that will be explained later. For every $i\in[s]$, we insert the piece $P_{i}=(v^{i}_{1}\ldots v^{i}_{\mu})$ between the vertices $u_{1}^{j_{i}},u_{2}^{j_{i}}$ and $u_{3}^{j_{i}},u_{4}^{j_{i}}$ of $D_{j_{i}}=(u_{1}^{j_{i}}u_{2}^{j_{i}}u_{3}^{j_{i}}u_{4}^{j_{i}})$ making up the square of a path $(u_{1}^{j_{i}}u_{2}^{j_{i}}v^{i}_{1}\ldots v^{i}_{\mu}u_{3}^{j_{i}}u_{4}^{j_{i}})$. We get a smoothed fragment $H^{\prime}$ with exactly the same number of edges as $H$ has. The crucial observation is that $H^{\prime}\cup\mathbf{W}$ still has a subgraph from $\mathcal{F}_{n}$. Moreover, this smoothed fragment satisfies the important ‘smoothness’ requirement: all powers of paths that it contains have length $o(\ln n)$.
@@
So, this modification of fragments allows to avoid problematic large closed subgraphs. However, it may reduce the total number of different fragments, that will eventually affect moments calculation. Let $\mathcal{H}_{n}=\left\{H(F),\,F\in\mathcal{F}_{n}(\overrightarrow{D})\right\}$ be the ($\mathbf{W}$-random) multiset of original $(F,\mathbf{W})$-fragments (we choose one fragment per $F$), and let $\mathcal{H}^{\prime}_{n}$ be the set of smoothed $(F,\mathbf{W})$-fragments. Although each $H\in\mathcal{H}_{n}$ maps to its smoothed version $H^{\prime}\in\mathcal{H}^{\prime}_{n}$, this correspondence $\varphi$ is not necessarily injective. Luckily, every pre-image is not large if we choose the rule of how $P_{i}$ are matched to $D_{j_{i}}$ randomly. We now explain how this matching is chosen. Consider a binomial random bipartite graph $\mathbf{B}$ with parts $V=[\chi]$ and $U=[n]$ and with edge probability $\beta:=n^{-1/3}/w$. Let $V^{\prime}\subset V$ be the set of all indices of diamonds that do not belong to a component of $H$ that contains some $P_{i}$. We know that $|V^{\prime}|\geq\frac{2}{3}|V|$. Then, we insert every $P_{i}=(v^{i}_{1}\ldots v^{i}_{\mu})$ into some $D_{j_{i}}$ so that
@@
  • there exists an edge between $j_{i}\in V$ and $\pi_{F}(v^{i}_{1})\in U$ in $\mathbf{B}$.
@@
Let $\Sigma$ be the set of all $(F,W,F^{\prime})$ such that $\mathbb{P}(\mathbf{F}=F,\mathbf{W}=W,\mathbf{F}^{\prime}=F^{\prime})>0$ and $|F\cap F^{\prime}|\leq\ell_{0}+5\chi$. For $(F,W,F^{\prime})\in\Sigma$ and a bipartite graph $B$ on $V\times U$, let us call the tuple $(B,F,W,F^{\prime})$ well-mixed, if $B[V^{\prime}\cup\{\pi_{F}(v^{i}_{1}),\,i\in[s]\}]$ has a matching that covers all vertices $\pi_{F}(v^{i}_{1}),\,i\in[s]$, where $v^{i}_{1}$ are the first vertices of the pieces of $H:=F\cap F^{\prime}$. For a fixed set $U^{\prime}\subset U$ of size exactly $|V^{\prime}|$, whp there exists a perfect matching between $V^{\prime}$ and $U^{\prime}$ in $\mathbf{B}[V^{\prime}\cup U^{\prime}]$ (see, e.g., [17, Theorem 4.1]). Therefore, recalling that the bound $|F\cap F^{\prime}|\leq\ell_{0}+5\chi$ implies that the number of pieces that we moved in $H$ is less than $2\chi/3\leq|V^{\prime}|$, we get that, for any fixed $(F,W,F^{\prime})\in\Sigma$, whp $(\mathbf{B},F,W,F^{\prime})$ is well-mixed. Also, due to Claim 6.2, Claim 6.4, and the described property of the distribution $\mathcal{Q}_{F,W}$, whp $(\mathbf{F},\mathbf{W},\mathbf{F}^{\prime})\in\Sigma$. Then,
@@
$\geq\sum_{(F,W,F^{\prime})\in\Sigma}\mathbb{P}((\mathbf{B},F,W,F^{\prime})\text{ is well-mixed}\mid\mathbf{F}=F,\mathbf{W}=W,\mathbf{F}^{\prime}=F^{\prime})$
@@
$=\sum_{(F,W,F^{\prime})\in\Sigma}\mathbb{P}((\mathbf{B},F,W,F^{\prime})\text{ is well-mixed})\times\mathbb{P}(\mathbf{F}=F,\mathbf{W}=W,\mathbf{F}^{\prime}=F^{\prime})$
$=(1-o(1))\mathbb{P}((\mathbf{F},\mathbf{W},\mathbf{F}^{\prime})\in\Sigma)=1-o(1),$
@@
Let $\Sigma$ be the (random) set of all pairs $\left(F\in\mathcal{F}_{n}(\overrightarrow{D}),W\in{{[n]\choose 2}\setminus(D_{1}\cup\ldots\cup D_{\chi})\choose m}\right)$ such that the random fragment $\mathbf{F}^{\prime}\sim\mathcal{Q}_{F,W}$ is nice. Due to the choice of $B$ and due to Claims 6.26.4, and 6.5, whp $(\mathbf{F},\mathbf{W})\in\Sigma$ (in the $(\mathbf{F},\mathbf{W},\mathbf{F}^{\prime})$-measure). Indeed,
@@
We now notice that the there exists a natural bijection between disjoint families $\mathcal{F}_{n}(\overrightarrow{D})$, $\mathcal{F}_{n}(\overrightarrow{D^{\prime}})$ for different tuples of diamonds $\overrightarrow{D}$ and $\overrightarrow{D^{\prime}}$. In particular, for $Y:=\sum_{\overrightarrow{D}}Y_{\overrightarrow{D}}$ we get that
@@ 6.3. Distribution of fragments
Let $\delta<1/2$ and let $\chi^{\prime}\leq n^{\delta}$ be a non-negative integer. Fix disjoint diamonds $D_{1},\ldots,D_{\chi+\chi^{\prime}}$ on $[n]$. Let $\overrightarrow{D}^{\prime}=(D_{1},\ldots,D_{\chi+\chi^{\prime}})$ be the concatenation of $\overrightarrow{D}$ with $(D_{\chi+1},\ldots,D_{\chi+\chi^{\prime}})$. We also fix some specific different positions $m_{1},\ldots,m_{\chi^{\prime}}\in[n]$ and denote $\overrightarrow{m}:=(m_{1},\ldots,m_{\chi^{\prime}})$. We then consider the family $\mathcal{F}^{m_{1},\ldots,m_{\chi^{\prime}}}_{n}(\overrightarrow{D}^{\prime})\subset\mathcal{F}_{n}(\overrightarrow{D})$ of all $F\in\mathcal{F}_{n}(\overrightarrow{D})$ such that, for every $i\in[\chi^{\prime}]$, the leftmost vertex of the diamond $D_{\chi+i}$ occupies the position $m_{i}$ in the ordering $\pi_{F}$ of $[n]$. For $W\in{{[n]\choose 2}\choose m}$, we call a $(\overrightarrow{D}^{\prime},\overrightarrow{m})$-grounded $(F,W)$-fragment any graph $F^{\prime}\cap F$ such that $F^{\prime}\in\mathcal{F}^{m_{1},\ldots,m_{\chi^{\prime}}}_{n}(\overrightarrow{D})$ and $F^{\prime}\subset F\cup W$. The proof of the following claim is identical to the proof of Claim 6.4; therefore, we omit it.
@@
Let $F\in\mathcal{F}^{m_{1},\ldots,m_{\chi^{\prime}}}_{n}(\overrightarrow{D}^{\prime})$ and let $\mathbf{W}\subset{[n]\choose 2}$ be a uniformly random $m$-subset. Then, with probability $1-O(n^{-1/2})$ in the $\mathbf{W}$-measure, we have that $|\mathbf{F}^{\prime}\cap F|\leq\ell_{0}+5\chi$ whp in the $\mathbf{F}^{\prime}$-measure, where $\mathbf{F}^{\prime}$ is a uniformly random $(\overrightarrow{D}^{\prime},\overrightarrow{m})$-grounded $(F,\mathbf{W})$-fragment.
We now fix $F\in\mathcal{F}_{n}(\overrightarrow{D})$, $W\in{{[n]\choose 2}\choose m}$, and define the distribution $\mathcal{Q}_{F,W}$. Let $\mathbf{F}^{\prime}$ be a uniformly random $(F,W)$-fragment. If $\mathbf{F}^{\prime}$ does not have closed subgraphs with at least $n/\ln n$ vertices, we let $\mathcal{Q}_{F,W}$ be the distribution of $\mathbf{F}^{\prime}$. Otherwise, set $F^{(1)}:=\mathbf{F}^{\prime}$. Let $C_{1},\ldots,C_{\chi^{\prime}}$ be the inclusion-maximal closed subgraphs of $F\cap F^{(1)}$ of length at least $n/\ln n$. Due to Claim 6.4, we have that $\chi^{\prime}=O(\log n)$ with probability $1-O(n^{-1/2})$ for (asymptotically) almost every pair $(F,W)$. If $\chi^{\prime}>n^{\delta}$, we let $\mathcal{Q}_{F,W}$ be the atomic distribution on $\{F\}$. If $\chi^{\prime}\leq n^{\delta}$, then, for every $i\in[\chi^{\prime}]$, let $D_{\chi+i}:=(u^{i}_{1}u^{i}_{2}u^{i}_{3}u^{i}_{4})$ be the diamond in $F$ that crosses $C_{i}$ in its leftmost vertices $u^{i}_{3},u^{i}_{4}$. We also let $m_{i}$ be the position of $u^{i}_{1}$ in the ordering $\pi_{F}$ of $[n]$. If there exists an $(\overrightarrow{D}^{\prime},\overrightarrow{m})$-grounded $(F,W)$-fragment $F^{\prime}$ such that $|F^{\prime}\cap F|\leq\ell_{0}+5\chi$, then we let $F^{(2)}$ be a uniformly random $(\overrightarrow{D}^{\prime},\overrightarrow{m})$-grounded $(F,W)$-fragment with this property. Otherwise, we let $\mathcal{Q}_{F,W}$ be the atomic distribution on $\{F\}$. Crucially, if $F^{(2)}$ is defined, then the two sets of all inclusion-maximal closed subgraphs of length at least $n/\ln n$ in $F\cap F^{(1)}$ and in $F\cap F^{(2)}$ do not overlap. Moreover, in this case, due to Claim 6.6, for at least $(1-O(1/\sqrt{n}))$-fraction of sets $W$, whp $|F^{(2)}\cap F|\leq\ell_{0}+5\chi$.
@@
We then proceed by induction. At step $2\leq j\leq n^{\delta}/\ln^{2}n$, we have that either $\mathcal{Q}_{F,W}$ has been defined, or graphs $F^{(1)},\ldots,F^{(j)}$ have been selected. If $j+1>n^{\delta}/\ln^{2}n$, then halt and let $\mathcal{Q}_{F,W}$ be the uniformly random distribution on $\{F^{(1)},\ldots,F^{(j)}\}$. Otherwise, let $C_{1},\ldots,C_{\chi^{\prime}}$ be the inclusion-maximal closed subgraphs of length at least $n/\ln n$ that belong to one of the fragments $F\cap F^{(1)},\ldots,F\cap F^{(j)}$. We assume by induction that $\chi^{\prime}=O(j\cdot\log n)$ for $(1-O(j/\sqrt{n}))$-fraction of sets $W$. If $\chi^{\prime}>n^{\delta}$, we let $\mathcal{Q}_{F,W}$ be the atomic distribution on $\{F\}$. If $\chi^{\prime}\leq n^{\delta}$, then, for every $i\in[\chi^{\prime}]$, let $D_{\chi+i}:=(u^{i}_{1}u^{i}_{2}u^{i}_{3}u^{i}_{4})$ be the diamond in $F$ that crosses $C_{i}$ in its leftmost vertices $u^{i}_{3},u^{i}_{4}$ and let $m_{i}$ be the position of $u^{i}_{1}$ in the ordering $\pi_{F}$ of $[n]$. If there exists an $(\overrightarrow{D}^{\prime},\overrightarrow{m})$-grounded $(F,W)$-fragment $F^{\prime}$ such that $|F^{\prime}\cap F|\leq\ell_{0}+5\chi$, then we let $F^{(j+1)}$ be a uniformly random $(\overrightarrow{D}^{\prime},\overrightarrow{m})$-grounded $(F,W)$-fragment $F^{\prime}$ with this property. Otherwise, we let $\mathcal{Q}_{F,W}$ be the atomic distribution on $\{F\}$. Due to Claim 6.6, for at least $(1-O(1/\sqrt{n}))$-fraction of sets $W$, whp $|F^{(j+1)}\cap F|\leq\ell_{0}+5\chi$.

### v3

@@ 6.2. Improving fragments
For a typical $F\in\mathcal{F}_{n}(\overrightarrow{D})$, let $H=H(F)=\mathbf{F}^{\prime}\cap F$ be an $(F,\mathbf{W})$-fragment, where $\mathbf{F}^{\prime}$ is sampled according to a certain distribution $\mathcal{Q}_{F,\mathbf{W}}$ over $\mathcal{F}_{n}(\overrightarrow{D})$ that will be defined later in Section 6.3. We will need this specific distribution, rather than the uniform distribution over all fragments, in order to prove that, on the last day, at least one fragment can be covered whp (see Section 6.7). In particular, this distribution has the following property: whp $|\mathbf{F^{\prime}}\cap F|\leq\ell_{0}+5\chi$, where $\mathbf{F}^{\prime}\sim\mathcal{Q}_{F,\mathbf{W}}$. In this section, we only need this particular property and do not need any other details about the distribution $\mathcal{Q}_{F,W}$. The definition of $\mathcal{Q}_{F,W}$ is quite cumbersome, thus we delay it, so that we do not interrupt the flow of this section. Therefore, we may assume that $H$ has size at most $\ell_{0}+5\chi$, for almost all pairs $(F,W)$.
@@
Recall that all graphs from $\mathcal{F}_{n}(\overrightarrow{D})$ contain every $D_{i}$, $i\in[\chi]$. Since $H=F\cap F^{\prime}$ for some $F^{\prime}\in\mathcal{F}_{n}(\overrightarrow{D})$, we get that every $D_{i}$ is a subgraph of $H$ as well. Moreover, since the distance in $F$ between any two $D_{i}$ is bigger than $2C\sqrt{n}$, we get that the connected components of $H$ that contain different $D_{i}$ are disjoint. Consider all inclusion-maximal closed subgraphs in $H$ with at least $10C\ln n/w$ vertices. For each such square of a path $v_{1}\ldots v_{h}$, we “cut” it into pieces of length $\mu:=\left\lfloor 10C\ln n/w\right\rfloor$ in the following way: $v_{1}v_{2}$; $v_{3}\ldots v_{2+\mu}$; $v_{3+\mu}\ldots v_{2+2\mu}$; $\ldots$. We let the last piece $v_{3+j\mu}v_{4+j\mu}\ldots$ to be of length at least 2 and at most $\mu+1$. From every such inclusion-maximal closed subgraph, remove all pieces (i.e. all the listed vertices as well as all the edges that contain at least one of them) but the first and the last one and glue the first and the last so that they make up the square of a path $(v_{1}v_{2}v_{3+j\mu}v_{4+j\mu}\ldots)$. Let $\mathcal{P}_{H}=\{P_{1},\ldots,P_{s}\}$ be the set of all pieces of length $\mu$ that we have removed ($P_{1},\ldots,P_{s}$ respect the order $\pi_{F}$ induced by $F$). Clearly, the number of pieces $s$ is less than $32\ell_{0}/\mu<\chi/$ 3, where $\chi$ is defined in (37).
@@
We then choose different $j_{1},\ldots,j_{s}\in[\chi]$, according to some rule, that will be explained later. In particular, any $D_{j_{i}}$ and any removed piece do not belong to the same connected component of the fragment. For every $i\in[s]$, we insert the piece $P_{i}=(v^{i}_{1}\ldots v^{i}_{\mu})$ between the first two vertices $u_{1}^{j_{i}},u_{2}^{j_{i}}$ and the last two vertices $u_{3}^{j_{i}},u_{4}^{j_{i}}$ of the diamond $D_{j_{i}}=(u_{1}^{j_{i}}u_{2}^{j_{i}}u_{3}^{j_{i}}u_{4}^{j_{i}})$ making up the square of a path $(u_{1}^{j_{i}}u_{2}^{j_{i}}v^{i}_{1}\ldots v^{i}_{\mu}u_{3}^{j_{i}}u_{4}^{j_{i}})$. We get a smoothed fragment $H^{\prime}$ with exactly the same number of edges as $H$ has (see Figure 1). The crucial observation is that $H^{\prime}\cup\mathbf{W}$ still has a subgraph from $\mathcal{F}_{n}$. Moreover, this smoothed fragment satisfies the important ‘smoothness’ requirement: all powers of paths that it contains have length $o(\ln n)$.
Figure 1. Construction of a smoothed fragment $H^{\prime}$: Connected components of $H$ and $H^{\prime}$ are represented by brown ovals; green squares indicate closed subgraphs with at least 6 vertices; blue squares represent pieces that were moved “inside” red diamonds.
@@
So, this modification of fragments allows to avoid problematic large closed subgraphs. However, it may reduce the total number of different fragments, that will eventually affect moments calculation. Let $\mathcal{H}_{n}=\left\{H(F),\,F\in\mathcal{F}_{n}(\overrightarrow{D})\right\}$ be the ($\mathbf{W}$-random) multiset of original $(F,\mathbf{W})$-fragments (we choose one fragment per almost every $F$) [^8] More precisely, the distribution of this multiset is induced by the product measure ×$\times_{F}\mathcal{Q}_{F,\mathbf{W}}$., and let $\mathcal{H}^{\prime}_{n}$ be the set of smoothed $(F,\mathbf{W})$-fragments. Although each $H\in\mathcal{H}_{n}$ maps to its smoothed version $H^{\prime}\in\mathcal{H}^{\prime}_{n}$, this correspondence $\varphi$ is not necessarily injective. Luckily, every pre-image is not large if we choose the rule of how $P_{i}$ are matched to $D_{j_{i}}$ randomly. We now explain how this matching is chosen. Consider a binomial random bipartite graph $\mathbf{B}$ with parts $V=[\chi]$ and $U=[n]$ and with edge probability $\beta:=n^{-1/3}/w$. Let $V^{\prime}\subset V$ be the set of all indices of diamonds that do not belong to a component of $H$ that contains some $P_{i}$. We know that $|V^{\prime}|\geq\frac{2}{3}|V|$. Then, we insert every $P_{i}=(v^{i}_{1}\ldots v^{i}_{\mu})$ into some $D_{j_{i}}$ so that
@@
  • there exists an edge between $j_{i}\in V$ and $\pi_{F}(v^{i}_{1})\in U$ in $\mathbf{B}$ (recall that $v^{i}_{1}$ is the first vertex of $P_{i}$).
@@
Let $\Gamma$ be the set of all $(F,W,F^{\prime})$ such that
  1. (1)
    $\mathbb{P}(\mathbf{F}=F,\mathbf{W}=W,\mathbf{F}^{\prime}=F^{\prime})>0$;
  2. (2)
    $|F\cap F^{\prime}|\leq\ell_{0}+5\chi$.
For $(F,W,F^{\prime})\in\Gamma$ and a bipartite graph $B$ on $V\times U$, let us call the tuple $(B,F,W,F^{\prime})$ well-mixed, if $B[V^{\prime}\cup\{\pi_{F}(v^{i}_{1}),\,i\in[s]\}]$ has a matching that covers all vertices $\pi_{F}(v^{i}_{1}),\,i\in[s]$, where $v^{i}_{1}$ are the first vertices of the pieces of $H:=F\cap F^{\prime}$. For a fixed set $U^{\prime}\subset U$ of size exactly $|V^{\prime}|$, whp there exists a perfect matching between $V^{\prime}$ and $U^{\prime}$ in $\mathbf{B}[V^{\prime}\cup U^{\prime}]$ (see, e.g., [18, Theorem 4.1]). Therefore, recalling that the bound $|F\cap F^{\prime}|\leq\ell_{0}+5\chi$ implies that the number of pieces that we moved in $H$ is less than $2\chi/3\leq|V^{\prime}|$, we get that, for any fixed $(F,W,F^{\prime})\in\Gamma$, whp $(\mathbf{B},F,W,F^{\prime})$ is well-mixed. Also, due to Claim 6.2, Claim 6.4, and the described property of the distribution $\mathcal{Q}_{F,W}$, whp $(\mathbf{F},\mathbf{W},\mathbf{F}^{\prime})\in\Gamma$. Then,
@@
$\geq\sum_{(F,W,F^{\prime})\in\Gamma}\mathbb{P}((\mathbf{B},F,W,F^{\prime})\text{ is well-mixed}\mid\mathbf{F}=F,\mathbf{W}=W,\mathbf{F}^{\prime}=F^{\prime})$
@@
$=\sum_{(F,W,F^{\prime})\in\Gamma}\mathbb{P}((\mathbf{B},F,W,F^{\prime})\text{ is well-mixed})\times\mathbb{P}(\mathbf{F}=F,\mathbf{W}=W,\mathbf{F}^{\prime}=F^{\prime})$
$=(1-o(1))\mathbb{P}((\mathbf{F},\mathbf{W},\mathbf{F}^{\prime})\in\Gamma)=1-o(1),$
@@
Let $\Sigma$ be the (random) set of all pairs $\left(F\in\mathcal{F}_{n}(\overrightarrow{D}),W\in{{[n]\choose 2}\setminus(D_{1}\cup\ldots\cup D_{\chi})\choose m}\right)$ such that the random fragment $\mathbf{F}^{\prime}\sim\mathcal{Q}_{F,W}$ is $(F,W)$-nice, where all fragments are chosen independently. Due to the choice of $B$ and due to Claims 6.26.4, and 6.5, whp $(\mathbf{F},\mathbf{W})\in\Sigma$ (in the $(\mathbf{F},\mathbf{W},\mathbf{F}^{\prime})$-measure). Indeed,
@@
We now notice that the there exists a natural bijection between disjoint families $\mathcal{F}_{n}(\overrightarrow{D})$, $\mathcal{F}_{n}(\overrightarrow{D^{\prime}})$ for different tuples of diamonds $\overrightarrow{D}$ and $\overrightarrow{D^{\prime}}$. In particular, for $Y:=\sum_{\overrightarrow{D}}Y_{\overrightarrow{D}}$ [^9] Recall that the random variable $Y_{\overrightarrow{D}}$ is defined in Section 6.1 as the number of $F\in\mathcal{F}_{n}(\overrightarrow{D})$ that have an $(F,\mathbf{W})$-fragment of size at most $\ell_{0}+5\chi$. we get that
@@ 6.3. Distribution of fragments
Let $\delta<1/2$ and let $\chi^{\prime}\leq n^{\delta}$ be a non-negative integer. Fix disjoint diamonds $D_{1},\ldots,D_{\chi+\chi^{\prime}}$ on $[n]$. Let $\overrightarrow{D}^{\prime}=(D_{1},\ldots,D_{\chi+\chi^{\prime}})$ be the concatenation of $\overrightarrow{D}=(D_{1},\ldots,D_{\chi})$ with $(D_{\chi+1},\ldots,D_{\chi+\chi^{\prime}})$, where $\overrightarrow{D}$ is as in Section 6.1. We also fix some specific different positions $m_{1},\ldots,m_{\chi^{\prime}}\in[n]$ and denote $\overrightarrow{m}:=(m_{1},\ldots,m_{\chi^{\prime}})$. We then consider the family $\mathcal{F}^{m_{1},\ldots,m_{\chi^{\prime}}}_{n}(\overrightarrow{D}^{\prime})\subset\mathcal{F}_{n}(\overrightarrow{D})$ of all $F\in\mathcal{F}_{n}(\overrightarrow{D})$ such that, for every $i\in[\chi^{\prime}]$,
  1. (1)
    the leftmost vertex of the diamond $D_{\chi+i}$ occupies the position $m_{i}$ in the ordering $\pi_{F}$ of $[n]$;
  2. (2)
    the diamond $D_{\chi+i}$ belongs to $F$ and the order of its vertices respects $\pi_{F}$.
For $W\in{{[n]\choose 2}\choose m}$, we call a $(\overrightarrow{D}^{\prime},\overrightarrow{m})$-grounded $(F,W)$-fragment any graph $F^{\prime}\cap F$ such that $F^{\prime}\in\mathcal{F}^{m_{1},\ldots,m_{\chi^{\prime}}}_{n}(\overrightarrow{D})$ and $F^{\prime}\subset F\cup W$. The proof of the following claim is identical to the proof of Claim 6.4; therefore, we omit it.
@@
Let $F\in\mathcal{F}^{m_{1},\ldots,m_{\chi^{\prime}}}_{n}(\overrightarrow{D}^{\prime})$ and let $\mathbf{W}\subset{[n]\choose 2}$ be a uniformly random $m$-subset. Then, with probability at least $1-\exp(-\Omega(\sqrt{n}))$ in the $\mathbf{W}$-measure, we have that $|\mathbf{F}^{\prime}\cap F|\leq\ell_{0}+5\chi$ whp in the $\mathbf{F}^{\prime}$-measure, where $\mathbf{F}^{\prime}$ is a uniformly random $(\overrightarrow{D}^{\prime},\overrightarrow{m})$-grounded $(F,\mathbf{W})$-fragment.
We now fix $F\in\mathcal{F}_{n}(\overrightarrow{D})$, $W\in{{[n]\choose 2}\choose m}$, and define explicitly a random fragment $\mathbf{F}^{*}=\mathbf{F}^{*}(F,W)$ that has the desired distribution $\mathcal{Q}_{F,W}$. Let $\mathbf{F}^{\prime}$ be a uniformly random subgraph of $F\cup W$ from $\mathcal{F}_{n}(\overrightarrow{D})$ satisfying $|\mathbf{F}^{\prime}\cap F|\leq\ell_{0}+5\chi$. If such a subgraph does not exist, we set $\mathbf{F}^{*}=F$. Otherwise, if $\mathbf{F}^{\prime}\cap F$ does not have closed subgraphs with at least $\sqrt{n}/\ln n$ vertices, we let $\mathbf{F}^{*}=\mathbf{F}^{\prime}$. If neither of the above two conditions is satisfied, set $\mathbf{F}^{(1)}:=\mathbf{F}^{\prime}$.
@@
We then proceed by induction. At step $1\leq j\leq n^{\delta}/\ln^{2}n$, we have that either $\mathbf{F}^{*}$ has been defined, or graphs $\mathbf{F}^{(1)},\ldots,\mathbf{F}^{(j)}$ have been sampled. If $j+1>n^{\delta}/\ln^{2}n$, then halt and let $\mathbf{F}^{*}$ be chosen uniformly at random from $\{\mathbf{F}^{(1)},\ldots,\mathbf{F}^{(j)}\}$. Otherwise, let $C_{1},\ldots,C_{\chi^{\prime}}$ be all the inclusion-maximal closed subgraphs of length at least $\sqrt{n}/\ln n$ that belong to one of the fragments $F\cap\mathbf{F}^{(1)},\ldots,F\cap\mathbf{F}^{(j)}$. For every $i\in[\chi^{\prime}]$, let $D_{\chi+i}:=(u^{i}_{1}u^{i}_{2}u^{i}_{3}u^{i}_{4})$ be the diamond in $F$ that crosses $C_{i}$ in its leftmost vertices $u^{i}_{3},u^{i}_{4}$ and let $m_{i}$ be the position of $u^{i}_{1}$ in the ordering $\pi_{F}$ of $[n]$. Resample $\mathbf{F}^{\prime}$ uniformly at random from $\mathcal{F}^{m_{1},\ldots,m_{\chi^{\prime}}}_{n}(\overrightarrow{D}^{\prime})$ under the condition that $|\mathbf{F}^{\prime}\cap F|\leq\ell_{0}+5\chi$. If such a graph does not exist, we set $\mathbf{F}^{*}=F$. Otherwise, if $\mathbf{F}^{\prime}\cap F$ does not have closed subgraphs with at least $\sqrt{n}/\ln n$ vertices, we let $\mathbf{F}^{*}=\mathbf{F}^{\prime}$. If neither of the above two conditions is satisfied, set $\mathbf{F}^{(j+1)}:=\mathbf{F}^{\prime}$.
Figure 2. Induction step in the process of construction of $\mathbf{F}^{*}$: Green squares represent inclusion-maximal closed subgraphs of length at least $\sqrt{n}/\ln n$; their leftmost vertices become rightmost vertices of the diamonds planted in $\mathbf{F}^{(j+1)}$. The order of vertices in the planted diamonds respects $\pi_{F}$ — in particular, grey edges of $F$ do not coincide with black edges of $\mathbf{F}^{(j)}$; the former grey edges are coloured red in $\mathbf{F}^{(j+1)}$.
This completes the construction of $\mathbf{F}^{*}$. Let us now prove that, in particular, its distribution satisfies the required property that we used in Section 6.2 — whp $|F\cap\mathbf{F}^{*}|\leq\ell_{0}+5\chi$. Let $\mathbf{W}\subset{[n]\choose 2}$ be a uniformly random $m$-subset. Due to Claim 6.4, with probability at least $1-\exp(-\Omega(\sqrt{n}))$, in the $\mathbf{W}$-measure, after the first step $j=1$, we either proceed to the next step or halt and output $\mathbf{F}^{*}$ that satisfies
  • P1
    $|\mathbf{F}^{*}\cap F|\leq\ell_{0}+5\chi$,
  • P2
    $\mathbf{F}^{*}$ does not have closed subgraphs with at least $\sqrt{n}/\ln n$ vertices.
Moreover, if we proceed to step $j=2$, then with probability at least $1-\exp(-\Omega(\sqrt{n}))$, in the $\mathbf{W}$-measure, the algorithm samples $\mathbf{F}^{(1)}$ that has $|\mathbf{F}^{(1)}\cap F|\leq\ell_{0}+5\chi$ (deterministically) and, therefore, the fragment $\mathbf{F}^{(1)}\cap F$ has less than $\ln^{2}n$ inclusion-maximal closed subgraphs of length at least $\sqrt{n}/\ln n$.
Then, for an arbitrary $1\leq j\leq n^{\delta}/\ln^{2}n$, assume that with probability at least $1-\exp(-\Omega(\sqrt{n}))$ in the $\mathbf{W}$-measure, at step $j$ either the algorithm halts and outputs $\mathbf{F}^{*}$ that satisfies P1 and P2, or it samples $\mathbf{F}^{(1)},\ldots,\mathbf{F}^{(j)}$ such that,
  • for every $i\in[j]$, $\mathbf{F}^{(i)}$ satisfies P1 and, therefore, $|F\cap\mathbf{F}^{(i)}|$ has less than $\ln^{2}n$ inclusion-maximal closed subgraphs of length at least $\sqrt{n}/\ln n$ and,
  • crucially, the $j$ sets of all inclusion-maximal closed subgraphs of length at least $\sqrt{n}/\ln n$ in $F\cap\mathbf{F}^{(1)},\ldots,F\cap\mathbf{F}^{(j)}$ do not overlap.
These sets identify the diamonds $C_{1},\ldots,C_{\chi^{\prime}}$, where $\chi^{\prime}<j\ln^{2}n<n^{\delta}$ whp. Due to Claim 6.6 and the union bound over the choice of $\chi^{\prime}$ diamonds, with probability at least $1-\exp(-\Omega(\sqrt{n}))$ in the $\mathbf{W}$-measure, we get that either the algorithm halts at step $j+1$ and outputs $\mathbf{F}^{*}$ that satisfies P1 and P2, or it samples at this step $\mathbf{F}^{(j+1)}$ that satisfies $|\mathbf{F}^{(j+1)}\cap F|\leq\ell_{0}+5\chi$ and proceeds to the next step.
We conclude that whp, in the $\mathbf{W}$-measure, the algorithm outputs $\mathbf{F}^{*}\sim\mathcal{Q}_{F,\mathbf{W}}$ that satisfies property P1, as required. Moreover, the algorithm is designed in such a way that whp, in the $\mathbf{W}$-measure,
  • P3